OpenStudy (anonymous):

solve the given inequality 6^6n<1/1296

OpenStudy (anonymous):

\[6^6n<\frac{1}{1296} \]\[n<\frac{1}{1296}*\frac{1}{6^6} \]\[n<\frac{1}{60466176} \]